# Parallel Point-to-Point Shortest Paths and Batch Queries

### v1

@@ Abstract.
We propose Orionet, efficient parallel implementations of Point-to-Point Shortest Paths (PPSP) queries using bidirectional search (BiDS) and other heuristics, with an additional focus on batch PPSP queries. We present a framework for parallel PPSP built on existing single-source shortest paths (SSSP) frameworks by incorporating pruning conditions. As a result, we develop efficient parallel PPSP algorithms based on early termination, bidirectional search, A search, and bidirectional A—all with simple and efficient implementations.
@@ 1. Introduction
We evaluated Orionet on both single and batch PPSP queries using 14 graphs with different types. For single PPSP queries, we tested Orionet using various distance percentile (1%, 50%, and 99% closest) query pairs. and compared it against two baselines, GraphIt (CGO’20) (Zhang et al., 2020) and MBQ (SPAA’24) (Zhang et al., 2024). Both of them only support parallel single PPSP and A using unidirectional search. In Sec. 6, we show our algorithm with bidirectional search achieves significant speedup over the baselines. On average, across all tests, our BiDS is 2.9$\times$ faster than GraphIt, and 6.8$\times$ faster than MBQ (even excluding MBQ’s timeout cases). Our bidirectional A is 4.4$\times$ and 6.2$\times$ faster than the A in GraphIt and MBQ, respectively. These results demonstrate the effectiveness of Orionet, as well as the proposed algorithmic insights in this paper. For batched queries, we provide in-depth experimental evaluation to compare the two solutions in Orionet (BiDS-based and SSSP-based) on various query graph patterns. Our code is publicly available (Dong et al., 2025). We believe this is the first systematic study of BiDS, A, and BiD-A for PPSP, and batch PPSP queries in the parallel setting.
@@ 2. Preliminaries
Computational Models.  We use the standard binary fork-join model (Cormen et al., 2009; Blelloch et al., 2020b). We assume a set of threads that share a common memory. A process can fork two child software threads to work in parallel. When both children complete, the parent process continues. We use the atomic operation write_min$(p,\mathit{v})$, which reads the memory location pointed to by $p$ and writes value $v$ to it if $v$ is smaller than the current value. It returns true if the update is successful and false otherwise.
@@ 6             ParallelForEach $v\in N(u)$ do
Stepping Algorithm Framework.  The stepping algorithm framework proposed in (Dong et al., 2021), is a high-level abstraction of many existing parallel SSSP algorithms, including Dijkstra (Dijkstra, 1959; Brodal et al., 1998), Bellman-Ford (Bellman, 1958; Ford Jr, 1956), $\Delta$-Stepping (Meyer and Sanders, 2003), $\rho$-Stepping (Dong et al., 2021), and more (Blelloch et al., 2016; Shi and Spencer, 1999). The framework (see Alg. 1) runs in steps and maintains a frontier $F$ as the set of vertices to be processed. In each step, it extracts vertices (Alg. 1) that have tentative distances within a certain threshold $\theta$ (Alg. 1) from the frontier and relaxes their neighbors in parallel (Alg. 1). If a vertex is successfully relaxed (Alg. 1), it will be added to the next frontier (Alg. 1). This framework applies to various algorithms by plugging in different functions for GetDist, which decides the threshold $\theta$ of the current step. The algorithmic ideas in this paper can be combined with many SSSP algorithms in the stepping framework. In Orionet, we use $\Delta^{*}$-stepping (Meyer and Sanders, 2003; Dong et al., 2021) as it is one of the most commonly used parallel SSSP algorithms. In $\Delta^{*}$-stepping, the $i$-th call of ${\mbox{\sc{GetDist}}}()$ returns $i\cdot\Delta$, where $\Delta$ is a user-defined parameter.
@@ 3. Algorithms for Point-to-Point Shortest Paths
In this section, we present our point-to-point shortest path (PPSP) framework and implementations of early termination (ET), A, BiDS, and BiD-A based on this framework. All these techniques have been shown to be effective for sequential PPSP. Although early termination and A have also been studied in parallel settings, we are aware of little work on using their bidirectional counterparts to achieve high-performance PPSP. A key challenge we observed is that the efficiency of sequential BiDS relies on strong ordering in vertex visits (Dijkstra’s order), which is not preserved in parallel SSSP algorithms. Parallel SSSP algorithms usually update vertices in batches or asynchronously, where strong ordering is not enforced. Therefore, in parallel, non-trivial adaptations are required to apply the ideas of BiDS, A, and BiD-A effectively to improve performance.
To overcome these challenges, we first design a PPSP framework adapted from the stepping algorithms framework in (Dong et al., 2021). Our framework eases the coding effort by abstracting the pruning conditions and, more importantly, facilitates our algorithm on batch PPSP queries. To incorporate bidirectional search, we design a pruning criterion independent of Dijkstra’s order and present its adaptation for BiD-A. Both our BiDS and BiD-A fit seamlessly into our framework and significantly improve the performance. In the following, we first present our framework (Sec. 3.1), then discuss early termination (Sec. 3.2) and A (Sec. 3.3) as examples to show the simplicity of our framework. We then present our algorithms on BiDS (Sec. 3.4) and BiD-A (Sec. 3.5) and prove the correctness.
@@ 1 $\mu\leftarrow+\infty,\delta[\cdot]\leftarrow+\infty$
2 ${\mbox{\sc{Init}}}()$ // Initialize the frontier $F$ and the distance $\delta[\cdot]$ for the sources
@@ 9                               $\mu={\mbox{\sc{UpdateDistance}}}(v^{\langle\oplus\rangle})$
The algorithm starts by initializing the frontier by function Init (Alg. 2) and then proceeds in steps, similar to a regular stepping algorithm. The main difference is that since searches can be from multiple sources (e.g., a bidirectional search or batch queries), each vertex may appear in the frontier multiple times (each from a different source). We use $u^{\langle\oplus\rangle}$ to denote an element in the frontier, where $u\in V$ is a vertex and $\oplus$ denotes the source of the search. For example, in bidirectional searches, a vertex $u$ can appear in both forward and backward searches. We use $u^{\langle+\rangle}$ and $u^{\langle-\rangle}$ to denote these two cases, respectively. As a result, the extracted vertices are those with the closest distances across multiple sources. If a vertex $u^{\langle\oplus\rangle}$ is extracted, we first determine whether it should be pruned using the function Prune (Alg. 2). If not, we process its neighbors regularly. During the algorithm, we maintain a global variable $\mu$, which is the current shortest distance from $s$ to $t$. If a vertex $v\in N(u)$ can be relaxed, we use UpdateDistance to update the current shortest distance $\mu$ with the path $s$-$v$-$t$, ensuring that $\mu$ eventually converges to the exact shortest distance from $s$ to $t$. If the neighbor $v$ does not satisfy the pruning condition, it is added to the frontier (Alg. 2).
@@
$\delta[t^{\langle-\rangle}]\leftarrow 0$, $F.{\mbox{\sc{Add}}}{}(t^{\langle-\rangle})$ For Prune$(v^{\langle-\rangle})$: $\delta[v^{\langle-\rangle}]+h_{B}(v)\geq\mu/2$
@@
  Multi-PPSP For Prune$(v^{\langle i\rangle})$: $\delta[v^{\langle i\rangle}]\geq\mu_{\max}[i]/2$ For Update$(v^{\langle i\rangle})$, let $q_{i}$ be the $i$-th vertex in $V_{q}$:
@@ 3.4. Bidirectional Search (BiDS)
Our BiDS.  We design our BiDS by plugging in the new Prune and UpdateDistance functions in Orionet, which are independent of Dijkstra’s order. Note that the key difference here is that, in parallel SSSP, an algorithm only knows the tentative distance on vertex $v$ but cannot determine whether and when it converges to the true distance. In contrast, in Dijkstra’s order, every vertex popped from the priority queue is settled with its true distance. Our BiDS is based on an observation that, in BiDS, given any upper bound $\mu$ of $d(s,t)$, we can skip a vertex $v$ with tentative distance $\delta[v]\geq\mu/2$. This indicates that any further searches beyond this point are wasteful and thus pruned. Unlike existing parallel PPSP work that does not apply or can only partially apply BiDS, our algorithm can apply BiDS throughout the entire search.
@@ 3.5. Bidirectional Astar (BiD-A)
With BiDS and A, it is natural to consider combining them as the bidirectional $A^{*}$ (BiD-A). However, BiD-A is challenging even in the sequential setting. Given two valid heuristic functions $h_{s}(v)$ estimating the distance to the source and $h_{t}(v)$ estimating the distance to the target, let $h_{F}(\cdot)$ be the heuristic used in the forward search (from $s$) and $h_{B}(\cdot)$ the heuristic in the backward search (from $t$). A simple approach is to use $h_{F}(\cdot)=h_{t}(\cdot)$ for the forward search and $h_{B}(\cdot)=h_{s}(\cdot)$ for the backward search and terminate when the two searches touch. Unfortunately, this solution is incorrect (Goldberg and Harrelson, 2005; Ikeda et al., 1994). Intuitively, as in A, we can consider the induced graph $G^{\prime}=(V,E,w^{\prime})$ with modified edge weights $w^{\prime}(u,v)=w(u,v)-h(u)+h(v)$. In the forward search, an edge $(u,v)$ has weight $w^{\prime}(u,v)=w(u,v)-h_{F}(u)+h_{F}(v)$, while its mirrored edge in the backward search has weight $w^{\prime}(v,u)=w(v,u)-h_{B}(v)+h_{B}(u)$, which are not equal. This difference breaks the consistency of the heuristic and thus leads to incorrect results in this simple solution. To fix this issue, the most commonly used technique (Goldberg and Harrelson, 2005) is to ensure $h_{B}(v)+h_{F}(v)=0$ for any vertex $v\in V$. Therefore, the modified edge weights $w^{\prime}$ are identical in both forward and backward searches. We will discuss how our solution achieves a consistent heuristic for BiDS below.
@@ 4.2. BiDS-Based Solutions
As discussed in Sec. 3, BiDS can greatly improve performance in a single PPSP query, so algorithmically, it is interesting to see how to integrate BiDS in the batch setting. One straightforward solution is to run BiDS for each query, which we refer to as the plain approach. However, a batch of queries often share common vertices. After revealing the common vertices, we can reduce the redundant searches originating from the same sources. More importantly, in BiDS, a target can also be considered as a source on undirected graphs. When each vertex can be either source or target and can be involved in multiple PPSP queries, we need to carefully redefine the conditions for pruning and updating the answers. We refer to this approach as the Multi-BiDS (or multi for short) approach. Our solution in Orionet is as follows. Since a vertex can be reached from multiple sources, for each vertex $u$ from the original graph, we maintain a copy of it from each of its neighbors in the query graph $G_{q}$. We use $u^{\langle i\rangle}$ to denote that vertex $u$ is reached from $q_{i}\in V_{q}$, and $\delta[u^{\langle i\rangle}]$ stores the tentative distance from $q_{i}$ to $u$ (i.e., $\delta[u^{\langle i\rangle}]=\delta_{q_{i}}[u]$). Our pruning idea is based on pruning the search from a vertex $q_{i}\in V_{q}$ when it passes the farthest midpoint of $q_{i}$ and $v\in N_{q}(q_{i})$.
@@
The space usage of this approach is $O(n\cdot|V_{q}|)$, which can be large for a very large query batch. In this case, we can simply process a subset of queries in turn, or use the SSSP-based solution mentioned below. We note that when the batch is very large, say $O(n)$, then some shortest-path algorithms with preprocessing (see Sec. 7) can be a better option.
@@ 6. Experiments
Table 3. Graphs information.$n$$=$ number of vertices. “$m$$=$ number of edges. “$D$$=$ diameter. “—LCC—%” $=$ ratio of the largest connected components.
@@ 6.1. Tested Algorithms
Baselines.  We compare Orionet with three SOTA baselines on single PPSP. The first is the original parallel $\Delta^{*}$-stepping SSSP algorithm from SPAA’21 (Dong et al., 2021), which is also the base implementation we build upon. The other two are GraphIt from CGO’20 (Zhang et al., 2020) and Multi-Bucket Queue (MBQ) from SPAA’24 (Zhang et al., 2024). Both of them only support unidirectional search with early termination (GI-ET and MBQ-ET) and A (GI-A and MBQ-A). Since MBQ does not support floating-point distances due to bitmasking, we round floating-point values to integers in its experiments, which favors MBQ over other implementations.
For batched queries, we did not find publicly available implementations for general batch PPSP, so we compare four different approaches in Orionet (all described in Sec. 4): 1) a plain BiDS-based solution running BiDS for all query pairs, 2) a BiDS-based solution using Multi-BiDS, 3) a plain solution running SSSP from all sources, and 4) an SSSP-based solution using VC.
@@ 6.3. Batch PPSP Queries
We present our experiments on batch PPSP queries in Fig. 7. We generate eight types of batch queries to simulate different use cases, as discussed in Sec. 1 and 4. Their query graphs are outlined in Fig. 7. The numbers in Fig. 7 represent relative running time normalized to the fastest approach on each graph and each query graph. To compare across queries, we fix the number of sources to six and vary only the queries between them. The tested algorithms are discussed in Sec. 6.1. Note that since the vertices are randomly picked, the expected distance is at about the 50-th percentile.
Overall, our Multi-BiDS achieves the best performance on about 70% of the tests. Even in the worst case (star), Multi-BiDS is within 50% slower than the fastest baseline on average. The speedups of Multi-BiDS over other baselines become larger when the query graph becomes denser (i.e., more sources are shared in the queries). For example, on cliques, Multi-BiDS is up to 5.41$\times$ faster (on CH5) even compared to the best baseline, and 2.06$\times$ on average. Similarly, on other query graph patterns where each vertex may be involved in more queries, such as chain, bipartite, and random, Multi-BiDS have an overwhelming advantage over the others.
Even on separate queries (three disjoint $s$-$t$ pairs) where no sources are shared among different queries, Multi-BiDS still benefits from running all queries in parallel. Compared to Plain-BiDS, Multi-BiDS achieves competitive performance on social and web graphs and achieves up to 84% speedups on road and $k$-NN graphs. This indicates another benefit of using Multi-BiDS on the batch: SSSP/PPSP algorithms usually suffer from insufficient parallelism because road and $k$-NN graphs are sparse and have large diameters. Batching the queries can naturally saturate the processors with sufficient jobs. On average, Multi-BiDS is 12% faster Plain-BiDS.
@@
More generally, we can observe that a VC-based SSSP solution is almost always better (up to 2.51$\times$ times faster) than a plain SSSP solution since the number of SSSPs running by VC is no more than the plain SSSP solution. As expected, on graphs where a small number of vertices can cover all edges, such as star and fork graphs, the VC-based solution can even outperform the solution based on Multi-BiDS, especially on social and web graphs. As mentioned, this is because SSSP on these graphs is highly optimized, which makes the additional work on maintaining BiDS more significant. Therefore, running one or two SSSPs may be cheaper than running BiDS from all six vertices.

### v2

@@ Abstract.
We propose Orionet, efficient parallel implementations of Point-to-Point Shortest Paths (PPSP) queries using bidirectional search (BiDS) and other heuristics, with an additional focus on batch PPSP queries. We present a framework for parallel PPSP built on existing single-source shortest paths (SSSP) frameworks by incorporating novel pruning conditions. As a result, we develop efficient parallel PPSP algorithms based on early termination, bidirectional search, A search, and bidirectional A—all with simple and efficient implementations.
@@ 1. Introduction
We evaluated Orionet on both single and batch PPSP queries using 14 graphs with different types. For single PPSP queries, we tested Orionet using various distance percentile (1%, 50%, and 99% closest) query pairs. and compared it against two baselines, GraphIt (CGO’20) (Zhang et al., 2020) and MBQ (SPAA’24) (Zhang et al., 2024). Both of them only support parallel single PPSP and A using unidirectional search. In Sec. 6, we show our algorithm with bidirectional search achieves significant speedup over the baselines. On average, across all tests, our BiDS is 2.9$\times$ faster than GraphIt, and 6.8$\times$ faster than MBQ (even excluding MBQ’s timeout cases). Our bidirectional A is 4.4$\times$ and 6.2$\times$ faster than the A in GraphIt and MBQ, respectively. These results demonstrate the effectiveness of Orionet, as well as the proposed algorithmic insights in this paper. For batched queries, we provide in-depth experimental evaluation to compare the two solutions in Orionet (BiDS-based and SSSP-based) on various query graph patterns. Our code is publicly available (Dong et al., 2025a). We believe this is the first systematic study of BiDS, A, and BiD-A for PPSP, and batch PPSP queries in the parallel setting.
@@ 2. Preliminaries
[…]
@@ 6     ParallelForEach $v\in N(u)$ do
[…]
@@ 3. Algorithms for Point-to-Point Shortest Paths
In this section, we present our point-to-point shortest path (PPSP) framework and implementations of early termination (ET), A, BiDS, and BiD-A based on this framework. All these techniques have been shown to be effective for sequential PPSP. Although early termination and A have also been studied in parallel settings, we are aware of little work on using their bidirectional counterparts to achieve high-performance PPSP. A key challenge we observed is that the efficiency of sequential BiDS relies on strong ordering in vertex visits (Dijkstra’s order), which is not preserved in parallel SSSP algorithms. Parallel SSSP algorithms usually update vertices in batches or asynchronously, where strong ordering is not enforced. Hence, in parallel, non-trivial adaptations are required to apply the ideas of BiDS, A, and BiD-A effectively to improve performance.
To overcome these challenges, we first design a PPSP framework adapted from the stepping algorithms framework in (Dong et al., 2021). Our framework eases the coding effort by abstracting the pruning conditions and, more importantly, facilitates our algorithm on batch PPSP queries. To incorporate bidirectional search, we design a pruning criterion independent of Dijkstra’s order and present its adaptation for BiD-A. Both our BiDS and BiD-A fit seamlessly into our framework and significantly improve the performance.
Note that both bidirectional search and A are considered heuristics, which do not improve the worst-case bounds. Our goal in this paper is to show that all these techniques can be the same effective in the parallel setting as in the classic sequential case. The work and span bounds for Orionet are upper-bounded by the underlying parallel SSSP algorithm (Dong et al., 2021).
In the rest of this section, we first present our framework (Sec. 3.1), then discuss early termination (Sec. 3.2) and A (Sec. 3.3) as examples to show the simplicity of our framework. We then present our new algorithms on BiDS (Sec. 3.4) and BiD-A (Sec. 3.5) and also prove the correctness.
@@ 1 $\mu\leftarrow+\infty,\delta[\cdot]\leftarrow+\infty$
[…]
@@ 9              $\mu={\mbox{\sc{UpdateDistance}}}(v^{\langle\oplus\rangle})$
The algorithm begins by initializing the frontier and setting the source distance to zero using the Init function (Alg. 2) and then proceeds in steps, similar to a regular stepping algorithm. The main difference is that since searches can be from multiple sources (e.g., a bidirectional search or batch queries), each vertex may appear in the frontier multiple times (each from a different source). We use $u^{\langle\oplus\rangle}$ to denote an element in the frontier, where $u\in V$ is a vertex and $\oplus$ denotes the source of the search. For example, in bidirectional searches, a vertex $u$ can appear in both forward and backward searches. We use $u^{\langle+\rangle}$ and $u^{\langle-\rangle}$ to denote these two cases, respectively. As a result, the extracted vertices are those with the closest distances across multiple sources. If a vertex $u^{\langle\oplus\rangle}$ is extracted, we first determine whether it should be pruned using the function Prune (Alg. 2). If not, we process its neighbors regularly. Note that whenever ${\mbox{\sc{Prune}}}(v)$ is called, $v$ satisfies $\delta[v]\neq\infty$ since $v$ is either the source or has been updated by other vertices. During the algorithm, we maintain a global variable $\mu$, which is the current shortest distance from $s$ to $t$. If a vertex $v\in N(u)$ can be relaxed, we use UpdateDistance to update the current shortest distance $\mu$ with the path $s$-$v$-$t$, ensuring that $\mu$ eventually converges to the exact shortest distance from $s$ to $t$. If the neighbor $v$ does not satisfy the pruning condition, it is added to the frontier (Alg. 2).
One may note that updating $\mu$ requires priority updates, but the contention here is light—existing analysis (Shun et al., 2013) shows that the content can generally be considered as logarithmic.
@@
[…]
@@
[…]
@@ 3.4. Bidirectional Search (BiDS)
Our BiDS.  We design our BiDS by plugging in the new Prune and UpdateDistance functions in Orionet, which are independent of Dijkstra’s order. Note that the key difference here is that, in parallel SSSP, an algorithm only knows the tentative distance on vertex $v$ but cannot determine whether and when it converges to the true distance. In contrast, in Dijkstra’s order, every vertex popped from the priority queue is settled with its true distance. Our BiDS is based on an observation that, in BiDS, given any upper bound $\mu$ of $d(s,t)$, we can skip a vertex $v$ with tentative distance $\delta[v]\geq\mu/2$. This indicates that any further searches beyond this point are wasteful and thus pruned. Unlike existing parallel PPSP work that does not apply or can only partially apply BiDS, our algorithm can apply BiDS throughout the entire search. We prove the correctness of our algorithm in  Thm. 3.3.
@@ 3.5. Bidirectional Astar (BiD-A)
[…]
@@
We note that the heuristic functions used in our paper are relatively straightforward. More advanced BiD-A heuristics have been developed, such as BAE (Sadhukhan, 2013) and DIBBS (Sewell and Jacobson, 2021). For a broader discussion on heuristic strategies in bidirectional search, we refer the reader to the survey (Sturtevant and Felner, 2018).
@@ 4.2. BiDS-Based Solutions
As discussed in Sec. 3, BiDS can greatly improve performance in a single PPSP query, so algorithmically, it is interesting to see how to integrate BiDS in the batch setting. A straightforward approach is to apply our parallel BiDS algorithm to each query one-by-one serially, which we refer to as the Plain approach. To further increase parallelism, we can instead run our parallel BiDS algorithm on all queries concurrently, which we refer to as the Plain approach. However, a batch of queries often share common vertices. After revealing the common vertices, we can reduce the redundant searches originating from the same sources. More importantly, in BiDS, a target can also be considered as a source on undirected graphs. When each vertex can be either source or target and can be involved in multiple PPSP queries, we need to carefully redefine the conditions for pruning and updating the answers. We refer to this approach as the Multi-BiDS (or multi for short) approach. Our solution in Orionet is as follows. Since a vertex can be reached from multiple sources, for each vertex $u$ from the original graph, we maintain a copy of it from each of its neighbors in the query graph $G_{q}$. We use $u^{\langle i\rangle}$ to denote that vertex $u$ is reached from $q_{i}\in V_{q}$, and $\delta[u^{\langle i\rangle}]$ stores the tentative distance from $q_{i}$ to $u$ (i.e., $\delta[u^{\langle i\rangle}]=\delta_{q_{i}}[u]$). Our pruning idea is based on pruning the search from a vertex $q_{i}\in V_{q}$ when it passes the farthest midpoint of $q_{i}$ and $v\in N_{q}(q_{i})$.
@@ 4.4. Takeaways and Further Discussions
Space Usage.  For BiDS-based approaches, the extra space needed for our Multi-BiDS algorithm is $O(n\cdot|V_{q}|)$, where $|V_{q}|$ is the number of vertices involved in the batch queries, instead of the number of queries in the batch. For Plain-BiDS, since we run the query one at a time, the extra space is $O(n)$. In comparison, Plain-BiDS requires $O(n\cdot|E_{q}|)$ extra space, as we need to maintain the distance array for each query. For SSSP-based approaches, the extra space is upper-bounded by $O(n\cdot|V_{q}|)$ for both algorithms, since the algorithm needs to instantiate the searches from at most $|V_{q}|$ sources.
Process Extremely Large Batches.  Orionet has a relatively high space usage if the given batches are very large. Therefore, for applications that require a huge number (e.g., $\Omega(n)$) of PPSP queries on the same input graph, some other methodologies with preprocessing may be a better choice. We discuss these ideas in Sec. 7.
@@ 6. Experiments
[…]
@@
Additional Experiments.  Due to page limits, we provide additional experimental results and analyses in the full version of the paper (Dong et al., 2025b), including comprehensive studies on distance percentiles, scalability, and memoization across all graphs, as well as experiments with larger batch sizes for batch PPSP queries.
@@ 6.1. Tested Algorithms
Baselines.  We compare Orionet with three SOTA baselines on single PPSP. The first is the original parallel $\Delta^{*}$-stepping SSSP algorithm from SPAA’21 (Dong et al., 2021), which is also the base implementation we build upon. The other two are GraphIt from CGO’20 (Zhang et al., 2020) and Multi-Bucket Queue (MBQ) from SPAA’24 (Zhang et al., 2024). Both of them contain PPSP algorithms with early termination (ET) and A in their source code and experiments in their papers. We directly port them (denoted as GI-ET/GI-A*/MBQ-ET/MBQ-A*) in our experiments. We select the best $\Delta$ for the two baselines as mentioned above, and all other settings and parameters remain unchanged from the given sample scripts. Since MBQ does not support floating-point distances due to bitmasking, we round floating-point values to integers in its experiments, which favors MBQ over other implementations.
For batched queries, we did not find publicly available implementations for general batch PPSP, so we compare four different approaches in Orionet (all described in Sec. 4): 1) Multi-BiDS: a BiDS-based solution using Multi-BiDS, 2) Plain-BiDS: a plain BiDS-based solution running parallel BiDS for each query one by one serially, 3) Plain-BiDS: another plain BiDS-based solution running parallel BiDS for all queries simultaneously, 4) VC-SSSP: an SSSP-based solution using VC, and 5) Plain-SSSP: a plain solution running parallel SSSP from all sources.
@@ 6.3. Batch PPSP Queries
We discuss the tested batch PPSP algorithms in Sec. 6.1 and present our experiments in Fig. 7. We generate eight types of batch queries to simulate different use cases, as discussed in Sec. 1 and 4. Their query graphs are outlined in Fig. 7. The numbers in Fig. 7 represent relative running time normalized to the fastest approach on each graph and each query graph. To compare across queries, we fix the number of sources to six and vary only the queries between them. The tested algorithms are discussed in Sec. 6.1. Note that since the vertices are randomly picked, the expected distance is at about the 50-th percentile.
Overall, our Multi-BiDS achieves the best performance on about 80% of the tests. Even in the worst case (star), Multi-BiDS is within 50% slower than the fastest baseline on average. The speedups of Multi-BiDS over other baselines become larger when the query graph becomes denser (i.e., more sources are shared in the queries). For example, on cliques, Multi-BiDS is up to 2.81$\times$ faster (on CH5) even compared to the best baseline, and 2.20$\times$ on average. Similarly, on other query graph patterns where each vertex may be involved in more queries, such as chain, bipartite, and random, Multi-BiDS have an overwhelming advantage over the others.
Even on separate queries (three disjoint $s$-$t$ pairs) where no sources are shared among different queries, Multi-BiDS still benefits from running all queries in parallel. Compared to Plain-BiDS, Multi-BiDS achieves competitive performance on social and web graphs and achieves up to 94% speedups on road and $k$-NN graphs. This indicates another benefit of using Multi-BiDS on the batch: SSSP/PPSP algorithms usually suffer from insufficient parallelism because road and $k$-NN graphs are sparse and have large diameters. Batching the queries can naturally saturate the processors with sufficient jobs. Compared to Plain-BiDS, Multi-BiDS achieves a 14% speedup on average, which we believe is due to better cache access patterns from collocating the queries. On average, Multi-BiDS is 25% faster than Plain-BiDS and 14% faster than Plain-BiDS.
@@
More generally, we can observe that a VC-based SSSP solution is almost always better (up to 2.55$\times$ times faster) than a plain SSSP solution since the number of SSSPs running by VC is no more than the plain SSSP solution. As expected, on graphs where a small number of vertices can cover all edges, such as star and fork graphs, the VC-based solution can even outperform the solution based on Multi-BiDS, especially on social and web graphs. As mentioned, this is because SSSP on these graphs is highly optimized, which makes the additional work on maintaining BiDS more significant. Therefore, running one or two SSSPs may be cheaper than running BiDS from all six vertices.